1 (base 6) + 5 (base 6)
4 (base 6) + 14 (base 6)
6543 (base 8) + 7426 (base 8)

Answer by stanbon(75887)

1 (base 6) + 5 (base 6) = 10(base 6)
-----------
4 (base 6) + 14 (base 6)
4+4 = 12(base 6)
1 + 1 = 2
Ans:: 22(base 6)
----------------------
6543 (base 8) + 7426 (base 8)
Adding from right to left::
3+6 = 1 carry 1
1 + 4 + 2 = 7
5 + 4 = 1 carry 1
1 + 6 + 7 = 6 carry 1
1 + 0 = 1
===
Final:: 16171(base 8)
------------
Cheers,
Stan H.
--------------